OpenStudy (anonymous):

Please help with me with this problem: Wite the expression in condensed form: 1/2logx -log y + log z I am getting log(square root of x / yz)

OpenStudy (anonymous):

z should be up top

OpenStudy (cwrw238):

1/2 log z = log z^(1/2) now use the other laws of logs log x - log y = = log x/y and log x + log y = log xy

OpenStudy (anonymous):

if it had been \(\frac{1}{2}\log(x)-\left(\log( y) + \log( z )\right)\) you would have been correct

OpenStudy (anonymous):

hope that is clear, it is like an "order of operations" question you are adding \(\log(z)\) not subtracting it so the \(z\) should be in the numerator, not the denominator just like \(10-4+3\) means first subtract 4, then add 3

OpenStudy (anonymous):

Thanks for the help

OpenStudy (anonymous):

yw